The sum of two numbers is 9 and their difference is 6. What are the two numbers?
Answer by nerdybill(7384)

The sum of two numbers is 9 antheir difference is 6. What are the two numbers?
Let x,y be the two number
then from "sum of two numbers is 9" we have:
x+y = 9 (equation 1)
.
and from "their difference is 6" we have:
x-y = 6 (equation 2)
.
applying "elimination method" we combine both equations together:
x+y = 9
x-y = 6
--------
2x = 15
x = 15/2
x = 7.5
.
to find y, we substitute above into equation 1:
x+y = 9
7.5+y = 9
y = 9-7.5
y = 1.5
.
the two numbers are 7.5 and 1.5